OpenStudy (aivantettet26):

At point (1,2) of the curve x^2-xy + y^2 = 3 find the equation of the tanget line

OpenStudy (aivantettet26):

tangent*

terenzreignz (terenzreignz):

You know tangent and derivative go together, right? Find \[\Large \frac{dy}{dx}\]

terenzreignz (terenzreignz):

It's not for anything that tough, just for finding the slope of the line. Now come, find the derivative... implicitly (as if there's any other way)

OpenStudy (aivantettet26):

2x+ y / x =y'?

terenzreignz (terenzreignz):

Whoa... I haven't even started yet... hang on XD

terenzreignz (terenzreignz):

\[\Large y^2-xy=3-x^2\]

OpenStudy (aivantettet26):

hahaha! XD no prob!

terenzreignz (terenzreignz):

\[\Large 2y \color{red}{y'}-x\color{red}{y'}-y= -2x\]

terenzreignz (terenzreignz):

\[\Large \color{red}{y'}(2y - x )=y-2x \]

terenzreignz (terenzreignz):

Did I make a mistake?

OpenStudy (aivantettet26):

but the answer is 2...n.n

OpenStudy (anonymous):

i see..is it only 2

terenzreignz (terenzreignz):

Wait what

OpenStudy (aivantettet26):

y = 2

terenzreignz (terenzreignz):

Right, you made an error in your differentiation :P

OpenStudy (aivantettet26):

xy is product rule right?

terenzreignz (terenzreignz):

This seems to hold: \[\Large \color{red}{y'}=\frac{2y - x }{y-2x}\]

terenzreignz (terenzreignz):

Yup. That's why differentiating it yields \[\Large x\color{red}{y'}+ y\]

OpenStudy (anonymous):

ya u're ryte...

terenzreignz (terenzreignz):

So, the only thing left to do is substitute the values (1,2) to the derivative (which happens to be the slope of the tangent line at that point. \[\Large \color{red}{y'}=\frac{2y - x }{y-2x}\]

OpenStudy (aivantettet26):

find the equation of the tangent line

terenzreignz (terenzreignz):

Find the slope. Then we'll do the rest. Just substitute x = 1 and y = 2 to \[\Large \color{red}{y'}=\frac{2y - x }{y-2x}\]

OpenStudy (aivantettet26):

4/3

terenzreignz (terenzreignz):

\[\Large \color{}{y'}=\frac{2\color{red}x - \color{blue}y }{\color{red}x-2\color{blue}y}\]

terenzreignz (terenzreignz):

Sorry, typo.

terenzreignz (terenzreignz):

Redo it, substitute x = 1 and y = 2 in \[\Large \color{}{y'}=\frac{2\color{red}x - \color{blue}y }{\color{red}x-2\color{blue}y}\]

OpenStudy (aivantettet26):

y' = 2(1) - 2 / 1-2(2)

terenzreignz (terenzreignz):

yes... go on :P

OpenStudy (aivantettet26):

0?

terenzreignz (terenzreignz):

That's right :) The slope is zero at that point. It can only mean we have a horizontal tangent line :P So... what is the equation of the HORIZONTAL LINE through the point (1,2) ?

OpenStudy (aivantettet26):

y = 2!

terenzreignz (terenzreignz):

Correcto :P

OpenStudy (anonymous):

so if use y=mx+c method 2=0(1)+c c =2 so y=2 Thanks terenzreignz coz just now i read the question wrongly :(

terenzreignz (terenzreignz):

You read it wrongly, I answered wrongly, we're even @Asha_Lvnr ^_^

OpenStudy (aivantettet26):

@Asha_Lvnr @terenzreignz thank you for your efforts!

OpenStudy (aivantettet26):

more questions to come!

OpenStudy (anonymous):

ahahaha u're welcome aivantettet26

OpenStudy (anonymous):

at last you solve it terenzreignz...hurrray